OpenStudy (anonymous):

what is the easiest way to find greatest common factors for large numbers, for example = gcf of 2541 and 5929

OpenStudy (asnaseer):

you need to express each number as a product of its unique primes, and then take the common factors from the two expressions.

OpenStudy (asnaseer):

in your example we would have:\[2451=3*7*11^{2}\]\[5929=7^{2}*11^{2}\]so c"common" factors here are 7 and 11^2 giving you the gcf of 7*11*11=847

OpenStudy (anonymous):

2541 = 3 * 7 * 11 * 11 5929 = 7 * 7 * 11 * 11 Common = 7 * 11 * 11 = 847 hence GCF = 847.

OpenStudy (anonymous):

Damn it! Haha nice answer asnaseer ;)

OpenStudy (asnaseer):

fewscrewsmissing - why "Damn it!"?

OpenStudy (anonymous):

Haha you beat me to the punch by like 3 seconds ;)

OpenStudy (asnaseer):

I didn't realise there was a race on. :-)